OpenStudy (anonymous):

(m^-1m^5 / m^-2)^-3 What is the simplified form of the expression?

OpenStudy (campbell_st):

well I'd apply the index laws inside the brackets 1st. and start with the numerator of the fraction. for multiplication, add the power \[x^a \times x^b = x^{a \times b}\] next you'll need the law for division, subtract the powers \[\frac{x^a}{x^b} = x^{a - b}\] then looking at the power of a power, multiply the powers \[(x^a)^b = x^{a \times b}\] I hope this helps you get the solution.

OpenStudy (anonymous):

thanks

OpenStudy (campbell_st):

its not a pretty question.... just asks you to use some laws...